Shift summary for day 07/01/99

Angela Bellavance (bellavan@rice.edu)
Thu, 1 Jul 1999 15:33:04 -0500 (CDT)

Shift Summary for day shift, 01/07/99

On Shift: NickS, AngelaB, ChristineP (UChicago summer student)

Hours of beam delivered at all intensities: 3.75
Hours of beam delivered at nominal intensity: 3.75
Hours of E832 data (special runs not included): 3.35
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:
Mounting new tapes.
TeV repairs.

========================================================================
Runs ended this shift:
Run 13777 - 13.3M events - BEAM_832
Run 13778 - 6.0M events - BEAM_832
Run 13779 - LASER_SCAN
Run 13780 - COSMICS (not saved)
Run 13781 - <no beam> - BEAM_832
Run 13782 - - BEAM_832

Tapes brought to FCC this shift:
QKK070-QKK072
QKC070-QKC072
QKS070-QKS072
QKB024

========================================================================
Accesses during shift:

Elizabeth and Valeri - fix HA channel 28 (replaced bad terminator on base)
and change voltages on VV'
Hogan - DC4 electronics work
Sydney - CsI channel 1136 problem shooting
Doug - Z position scan of chambers
BobT - restabalize abort gap signal and adjust HV on MU2
Nick and Christine - tour

========================================================================
Overview of shift:
Some E832 data-taking, mostly TeV downtime.
CsI laser scan and accesses made during beam downtime.
HA channel 28 fixed.
Voltages on VV' adjusted.
Z-position laser scan done.
CsI channel #1136 identified as having bad ground connection.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13777 7:20 Tapes fill up, run ends.
13778 1 7:37 Do fresh_start and begin with new tapes.
7:40 Notice livetime is low. Call MCR. They briefly
reduce intensity and reboot QXR.
Mount an on-split tape (COSMICS).
8:20 Nick vaults tapes.
8:40 MCR calls to check on our beam quality. They said
any further improvements might be
beam-destructive. We opt to continue as-is.
8:56 Booster radiation detector repair (beam gone).
Est. 5 minutes.
9:08 Beam back to users.
9:32 Mount two more on-split tapes.
10:11 MCR calls. Expect 1 hour downtime at 11:00.
10:20 Downtime estimate extended to 2.5 hours.
Appropriate people paged.
10:53 KTEVANA errors. Running KFREND fixes it.
11:00 Beam goes away for TeV magnet repairs. Run ended.
13779 1 11:12 Laser scan started.
11:15 People go on access: Elizabeth, Hogan, Valeri,
Doug, BobT. Later Nick, Christine. Later Sydney.
11:56 Laser scan ends. JasonH goes to swap a DYC.
13780 - 12:06 JasonH runs cosmics to test DYC swap.
13:12 Message: "Turning on MI and TeV. Est. 30 minutes."
13:15 Everyone out of access. New run initialized.
13781 1 13:20 Run started - no beam yet.
14:02 Message: "TeV hops problem, investigating."
What are they doing, brewing beer?!?
14:40 No beam yet, call MCR. Initial problem fixed, but
working on kicker magnet. Est. 15-20 minutes.
14:50 Run stopped to load new HV values (Sydney &
JasonH).
- 14:54 TLD location rad. monitors are replaced.
13782 1 15:00 Gas checklist done. New run started.
15:30 Beam is trying to come back, but aborting. Hand
off to Peter and Sasha.

========================================================================
Follow up Notes on Previous Shift:

Laser scan done, PeterSh notified.

========================================================================
Items to Follow up on Next Shift:

CsI channel #1136 needs to be fixed (bad dynode / ground connection) next
time we have > 1 hr. beam loss.

Regards, Angela

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke